A woman has a daughter and a son. The son is 3 years older than the daughter. In one year the woman will be six times as old as the daughter is now, and in 10 years, she will be 14 years older than the combined ages of her children. Whats the womans present age?
Answer by user_dude2008(1862)

s=d+3 ----> d=s-3
w+1=6d ----> w=6d-1
w+10=s+d+3*14


w+10=s+d+42
6d-1+10=s+d+42
6(s-3)-1+10=s+s-3+42
6s-18+9=2s+39
6s-9=2s+39
6s-2s=9+39
4s=48
s=12
d=s-3=12-3=9
d=9
w=6d-1=6(9)-1=54-1=53

Woman 53 yrs old
Son 12 yrs old
Daughter 9 yrs old